Suppose you wanted to recreate galileos experiment so you go out with a friend into a fueld at night and stand 10,000 m away from each other. you then light a flashlight, how long will it take for the flash of light to reach your friends eyes?
Answer #1

We know that speed of light, c = 3*10^8 m/s

distance between you and your friend = d = 10000 m

Now we know that:

time = distance/Velocity

t = d/c

t = 10000/(3*10^8) = 3.33*10^-5 sec

(If need answer in one significant figure, then approximate t = 3*10^-5 sec)

Let me know if you've any query.